# An optimal parallel-in-time preconditioner for parabolic optimal control problems

### v1

@@ 3 Numerical Result
$\gamma$ $h$ DoF GMRES-$\mathcal{P}_{\epsilon}$
Iter CPU $e_{h}$
$10^{-10}$ $2^{-5}$ 61504 4 0.185 1.03e-3
$2^{-6}$ 508032 4 0.775 5.17e-4
$2^{-7}$ 4129024 4 5.805 2.59e-4
$2^{-8}$ 33292800 4 54.357 1.30e-4
$10^{-8}$ $2^{-5}$ 61504 6 0.222 1.03e-3
$2^{-6}$ 508032 6 1.059 5.17e-4
$2^{-7}$ 4129024 6 8.241 2.59e-4
$2^{-8}$ 33292800 7 119.212 1.30e-4
$10^{-6}$ $2^{-5}$ 61504 8 0.339 1.02e-3
$2^{-6}$ 508032 10 2.046 5.15e-4
$2^{-7}$ 4129024 11 18.462 2.57e-4
$2^{-8}$ 33292800 13 194.312 1.29e-4
$10^{-4}$ $2^{-5}$ 61504 14 0.668 9.82e-4
$2^{-6}$ 508032 15 2.982 4.92e-4
$2^{-7}$ 4129024 13 24.479 2.46e-4
$2^{-8}$ 33292800 11 199.911 1.23e-4
$10^{-2}$ $2^{-5}$ 61504 11 0.516 4.03e-3
$2^{-6}$ 508032 9 1.973 2.17e-3
$2^{-7}$ 4129024 8 16.740 1.13e-3
$2^{-8}$ 33292800 7 146.326 5.76e-4
$1$ $2^{-5}$ 61504 6 0.337 2.85e-2
$2^{-6}$ 508032 6 1.610 1.43e-2
$2^{-7}$ 4129024 6 13.809 7.20e-3
$2^{-8}$ 33292800 6 131.204 3.61e-3
@@ 4 Conclusions
In this study, we developed a novel PinT preconditioner for all-at-once linear systems from optimal control problems with parabolic equations, using $\epsilon$-circulant matrices which can be efficiently implemented by fast Fourier transforms. Our approach enhances the convergence rate of the GMRES method, maintaining linearity with a theoretically estimated $\epsilon$ and independence from matrix size and the regularization parameter. Numerical experiments confirm the effectiveness and efficiency of our preconditioner.
@@ Acknowledgments
The work of Sean Y. Hon was supported in part by the Hong Kong RGC under grant 22300921 and a start-up grant from the Croucher Foundation. The work of Xue-lei Lin was partially supported by research grants: 12301480 from NSFC, HA45001143 from Harbin Institute of Technology, Shenzhen, HA11409084 from Shenzhen.

### v2

@@ 3 Numerical Result
Table 6: Results of GMRES-$\mathcal{P}_{\epsilon}$ for Example 3.2 with $\gamma=10^{-4},10^{-2},1$
$\gamma$ $h$ $\tau$ DoF GMRES-$\mathcal{P}_{\epsilon}$
Iter CPU $e_{h,\tau}$
$10^{-4}$ $2^{-6}$ $2^{-6}$ 508032 20 1.12 3.84e-06
$2^{-7}$ 1016064 18 1.92 1.92e-06
$2^{-8}$ 2032128 15 3.12 9.59e-07
$2^{-7}$ $2^{-6}$ 2064512 20 4.46 3.84e-06
$2^{-7}$ 4129024 18 7.60 1.92e-06
$2^{-8}$ 8258048 15 11.57 9.59e-07
$2^{-8}$ $2^{-6}$ 8323200 20 16.27 3.84e-06
$2^{-7}$ 16646400 18 29.74 1.92e-06
$2^{-8}$ 33292800 15 49.14 9.59e-07
$10^{-2}$ $2^{-6}$ $2^{-6}$ 508032 12 0.61 3.50e-04
$2^{-7}$ 1016064 11 1.19 1.75e-04
$2^{-8}$ 2032128 10 2.13 8.75e-05
$2^{-7}$ $2^{-6}$ 2064512 12 2.52 3.50e-04
$2^{-7}$ 4129024 11 4.78 1.75e-04
$2^{-8}$ 8258048 10 8.01 8.75e-05
$2^{-8}$ $2^{-6}$ 8323200 12 9.57 3.50e-04
$2^{-7}$ 16646400 11 18.97 1.75e-04
$2^{-8}$ 33292800 10 33.68 8.75e-05
$1$ $2^{-6}$ $2^{-6}$ 508032 10 0.52 1.85e-02
$2^{-7}$ 1016064 8 0.94 9.24e-03
$2^{-8}$ 2032128 8 1.87 4.61e-03
$2^{-7}$ $2^{-6}$ 2064512 10 2.28 1.85e-02
$2^{-7}$ 4129024 8 3.75 9.24e-03
$2^{-8}$ 8258048 8 7.23 4.61e-03
$2^{-8}$ $2^{-6}$ 8323200 10 8.16 1.85e-02
$2^{-7}$ 16646400 8 14.61 9.24e-03
$2^{-8}$ 33292800 8 29.12 4.61e-03
Table 7: Results of GMRES-$\mathcal{P}_{\epsilon}$ for Example 3.2 with various regularization parameters $\gamma$
$\gamma$ $h$ $\tau$ DoF GMRES-$\mathcal{P}_{\epsilon}$
Iter CPU $e_{h,\tau}$
$10^{-10}$ $2^{-5}$ $2^{-5}$ 61504 6 0.07 6.05e-06
$2^{-6}$ 123008 6 0.12 1.50e-06
$2^{-7}$ 246016 6 0.24 3.75e-07
$2^{-8}$ 492032 6 0.36 9.37e-08
$2^{-9}$ 984064 6 0.65 2.34e-08
$2^{-10}$ 1968128 8 1.66 5.85e-09
$2^{-11}$ 3936256 8 3.02 4.93e-09
$10^{-8}$ $2^{-5}$ $2^{-5}$ 61504 6 0.06 6.05e-06
$2^{-6}$ 123008 8 0.15 1.50e-06
$2^{-7}$ 246016 8 0.29 3.75e-07
$2^{-8}$ 492032 10 0.61 9.37e-08
$2^{-9}$ 984064 11 1.21 2.35e-08
$2^{-10}$ 1968128 13 2.83 5.91e-09
$2^{-11}$ 3936256 16 6.03 1.66e-09
$10^{-6}$ $2^{-5}$ $2^{-5}$ 61504 10 0.09 6.05e-06
$2^{-6}$ 123008 12 0.20 1.51e-06
$2^{-7}$ 246016 15 0.50 3.81e-07
$2^{-8}$ 492032 18 1.03 9.94e-08
$2^{-9}$ 984064 19 1.92 2.84e-08
$2^{-10}$ 1968128 17 3.46 9.56e-09
$2^{-11}$ 3936256 15 5.62 3.71e-09
$10^{-4}$ $2^{-5}$ $2^{-5}$ 61504 19 0.17 7.67e-06
$2^{-6}$ 123008 20 0.35 3.84e-06
$2^{-7}$ 246016 18 0.63 1.92e-06
$2^{-8}$ 492032 15 0.89 9.59e-07
$2^{-9}$ 984064 13 1.29 4.79e-07
$2^{-10}$ 1968128 11 2.15 2.40e-07
$2^{-11}$ 3936256 9 3.76 1.20e-07
$10^{-2}$ $2^{-5}$ $2^{-5}$ 61504 13 0.11 7.00e-04
$2^{-6}$ 123008 12 0.22 3.50e-04
$2^{-7}$ 246016 11 0.39 1.75e-04
$2^{-8}$ 492032 10 0.55 8.75e-05
$2^{-9}$ 984064 10 0.97 4.38e-05
$2^{-10}$ 1968128 8 1.72 2.19e-05
$2^{-11}$ 3936256 8 3.25 1.09e-05
$1$ $2^{-5}$ $2^{-5}$ 61504 10 0.12 3.72e-02
$2^{-6}$ 123008 10 0.20 1.85e-02
$2^{-7}$ 246016 8 0.30 9.24e-03
$2^{-8}$ 492032 8 0.50 4.61e-03
$2^{-9}$ 984064 8 0.78 2.31e-03
$2^{-10}$ 1968128 8 1.54 1.15e-03
$2^{-11}$ 3936256 8 3.18 5.76e-04
Table 8: Results of GMRES-$\mathcal{P}$ for Example 3.2 with various regularization parameters $\gamma$
$\gamma$ $h$ $\tau$ DoF GMRES-$\mathcal{P}$
Iter CPU $e_{h,\tau}$
$10^{-10}$ $2^{-5}$ $2^{-5}$ 61504 6 0.12 6.00e-06
$2^{-6}$ 123008 6 0.15 1.48e-06
$2^{-7}$ 246016 6 0.21 3.64e-07
$2^{-8}$ 492032 6 0.36 8.85e-08
$2^{-9}$ 984064 6 0.69 2.12e-08
$2^{-10}$ 1968128 8 1.73 5.19e-09
$2^{-11}$ 3936256 8 3.43 4.93e-09
$10^{-8}$ $2^{-5}$ $2^{-5}$ 61504 6 0.06 6.00e-06
$2^{-6}$ 123008 8 0.15 1.48e-06
$2^{-7}$ 246016 8 0.26 3.64e-07
$2^{-8}$ 492032 10 0.62 8.86e-08
$2^{-9}$ 984064 11 1.35 2.13e-08
$2^{-10}$ 1968128 13 3.09 5.24e-09
$2^{-11}$ 3936256 16 6.79 1.66e-09
$10^{-6}$ $2^{-5}$ $2^{-5}$ 61504 10 0.09 6.01e-06
$2^{-6}$ 123008 12 0.25 1.49e-06
$2^{-7}$ 246016 15 0.55 3.70e-07
$2^{-8}$ 492032 18 1.28 9.42e-08
$2^{-9}$ 984064 19 2.38 2.61e-08
$2^{-10}$ 1968128 17 4.08 8.56e-09
$2^{-11}$ 3936256 15 6.59 3.28e-09
$10^{-4}$ $2^{-5}$ $2^{-5}$ 61504 19 0.28 7.67e-06
$2^{-6}$ 123008 20 0.59 3.84e-06
$2^{-7}$ 246016 18 0.76 1.92e-06
$2^{-8}$ 492032 15 1.06 9.59e-07
$2^{-9}$ 984064 13 1.58 4.79e-07
$2^{-10}$ 1968128 11 2.68 2.40e-07
$2^{-11}$ 3936256 9 4.56 1.20e-07
$10^{-2}$ $2^{-5}$ $2^{-5}$ 61504 13 0.17 7.00e-04
$2^{-6}$ 123008 12 0.25 3.50e-04
$2^{-7}$ 246016 11 0.45 1.75e-04
$2^{-8}$ 492032 10 0.67 8.75e-05
$2^{-9}$ 984064 10 1.21 4.38e-05
$2^{-10}$ 1968128 8 2.01 2.19e-05
$2^{-11}$ 3936256 8 3.60 1.09e-05
$1$ $2^{-5}$ $2^{-5}$ 61504 10 0.12 3.72e-02
$2^{-6}$ 123008 10 0.23 1.85e-02
$2^{-7}$ 246016 10 0.40 9.24e-03
$2^{-8}$ 492032 10 0.72 4.61e-03
$2^{-9}$ 984064 10 1.28 2.31e-03
$2^{-10}$ 1968128 10 2.43 1.15e-03
$2^{-11}$ 3936256 10 4.80 5.76e-04
Figure 2: Convergence factor per iteration for Example 3.2 with various regularization parameters $\gamma$.
@@ 4 Conclusions
In this study, we developed a novel PinT preconditioner for the all-at-once linear systems from optimal control problems with parabolic equations, using $\epsilon$-circulant matrices which can be efficiently implemented by fast Fourier transforms. Our approach enhances the convergence rate of the GMRES method, maintaining linearity with a theoretically estimated $\epsilon$, and remains independent of both the matrix size and the regularization parameter. Numerical experiments confirm the effectiveness and efficiency of our preconditioner. As part of our future work, we intend to generalize our preconditioning strategies to accommodate more complex problem settings—particularly those involving parabolic optimal control problems discretized in time using the second-order Crank–Nicolson scheme, as well as optimal control problems constrained by fractional time-dependent diffusive equations.
@@ Acknowledgments
The work of Sean Y. Hon was supported in part by NSFC under grant 12401544, the Hong Kong RGC under grant 22300921 and a start-up grant from the Croucher Foundation. The work of Xue-Lei Lin was partially supported by research grants: 12301480 from NSFC, HA45001143 from Harbin Institute of Technology, Shenzhen, HA11409084 from Shenzhen.